What equal charges should be placed on earth and moon to neutralize their gravitation attraction

Grade:12

2 Answers

Md hasan
5 years ago
1Secondary SchoolScience 13 pointsWhat equal charges would have to be placed on earth and moon to neutralize their gravitaional attraction?given, mass of earth =1025 kg ,?AdvertisementAsk for details FollowReport by Jawaidsaman5134 20.06.2017AnswersAlbert01Albert01AceTo neutralize their gravity, the force of repulsion must equal the force of attraction. Universal Gravitational force = (G * m1 * m2) ÷ d^2 G = 6.67 * 10^-11 m1 = mass of earth = 5.98 * 10^24 kg m2 = mass of moon = 7.36 * 10^22 kg d = distance from earth to moon Force caused by charges = (k * q1 * q2) ÷ d^2 k = 9 * 10^9 q1 = charge on earth q2 = charge on moon d = distance from earth to moon Force caused by charges = Universal Gravitational force (k * q1 * q2) ÷ d^2 = (G * m1 * m2) ÷ d^2 Multiply both sides by d^2 (k * q1 * q2) = (G * m1 * m2) Divide both sides by k q1 * q2 = (G * m1 * m2) ÷ k The charges are equal, so q1 * q2 = q^2 q^2 = (G * m1 * m2) ÷ k q = √ (G * m1 * m2 ÷ k) q = √ (6.67 * 10^-11 * 5.98 * 10^24 * 7.36 * 10^22 ÷ 9 * 10^9) q ≈ 3.26 * 10^27 Coulombs
Md hasan
5 years ago
To neutralize their gravity, the force of repulsion must equal the force of attraction. Universal Gravitational force = (G * m1 * m2) ÷ d^2 G = 6.67 * 10^-11 m1 = mass of earth = 5.98 * 10^24 kg m2 = mass of moon = 7.36 * 10^22 kg d = distance from earth to moon Force caused by charges = (k * q1 * q2) ÷ d^2 k = 9 * 10^9 q1 = charge on earth q2 = charge on moon d = distance from earth to moon Force caused by charges = Universal Gravitational force (k * q1 * q2) ÷ d^2 = (G * m1 * m2) ÷ d^2 Multiply both sides by d^2 (k * q1 * q2) = (G * m1 * m2) Divide both sides by k q1 * q2 = (G * m1 * m2) ÷ k The charges are equal, so q1 * q2 = q^2 q^2 = (G * m1 * m2) ÷ k q = √ (G * m1 * m2 ÷ k) q = √ (6.67 * 10^-11 * 5.98 * 10^24 * 7.36 * 10^22 ÷ 9 * 10^9) q ≈ 3.26 * 10^27 Coulombs